let i=<x,2> and=<x,3> be ideal in z[x] .prove that IJ=<x,6>


1
Expert's answer
2022-03-07T17:23:02-0500

Let us prove it by double inclusion.

First of all, "3x\in IJ, 2x\in IJ \text{ so } 3x-2x=x\in IJ". Also, "3\cdot 2 = 6\in IJ", so we have "<x, 6>\subseteq IJ".

Now, let us prove the other inclusion. An element of "I" is of a form "x\cdot P(x) + 2a_0" and an element of "J" is of a form "x \cdot Q(x) + 3b_0". Therefore, an element of "IJ" is of a form "x\cdot (xP(x)Q(x)+2a_0 Q(x)+3b_0P(x)) + 6a_0b_0". We see, that it is an element of "<x,6>", so we conclude that

"IJ = <x,6>"